OpenStudy (anonymous):

The band at Crest Elementary School has 9 flutes, 6 clarinets, 3 trumpets, 2 trombones, 1 tuba, and 3 French horns. If one instrumentalist is chosen from the band, what is the probability that the student plays the flute?

OpenStudy (goformit100):



OpenStudy (anonymous):

9/total possibilities 9/(9+6+3+2+1+3) =9/24 = 3/8

OpenStudy (goformit100):

yes